OpenStudy (anonymous):

Write as a single log: (log3^x + log3^12) - log3^2

OpenStudy (allank):

Hint: logA+logB-logC = log((A*B)/C)

OpenStudy (cwrw238):

use laws of logs: log a + log b = log ab log a - log b = log a/b log^(na) = n log a

OpenStudy (allank):

True. Just a bit of adjustment, right @cwrw238? log a^n = n log a

OpenStudy (cwrw238):

yea - sorry log a^n = n log a

OpenStudy (anonymous):

so it is log(3x*3x^12)(3^2)

OpenStudy (allank):

Hmm, not quite. log3^x + log3^12- log3^2 Combining the first two, log(3^x*3^12)-log3^2 Then combining those gives: log((3^x*3^12)/3^2)

OpenStudy (anonymous):

 =(x+12)/2